OpenStudy (anonymous):

factor this expression 9r(squared) -4s(squared) divided by 9r+6s

OpenStudy (dumbcow):

factor top and bottom independently top is of the form (a^2-b^2) = (a-b)(a+b) a^2 = 9r^2 --> a = 3r b^2 = 4s^2 --> b = 2s = (3r-2s)(3r+2s) bottom has common factor of 3 =3(3r+2s) cancel (3r+2s) on top and bottom leaving (3r-2s)/3